OpenStudy (anonymous):

How do you solve the equation 4^(9n+4)=256

OpenStudy (kinggeorge):

You would take the log base 4 of both sides.\[\large \log_4(4^{9n+4})=\log_4(256)\]\[9n+4=\log_4(4^4)\]\[9n+4=4\]So n=0

OpenStudy (dream95):

use logarithm rules to solve it